The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu, Dominik J. Bomans, and Eva K. Grebel

                                  mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                 http://www.astro.uni-wuerzburg.de/~grebel/MCNews.html

No. 7 October 21, 1996

Table of Contents

News and Views..............1 Paper Abstracts..............2 Thesis Abstract...............6 Job Advertisement..........7

News and Views

IAU Symposium on the Magellanic Clouds

As announced in the last MC Newsletter, an IAU Symposium on the Magellanic Clouds is being planned. We are pleased with the responses from the community, and we thank those who have volunteered to help. A Scientific Organization Committee representing a diversity of disciplines and countries is being assembled. A formal proposal for the symposium will be forthcoming soon. Suggestions on the topics and format of the program are still welcome. Updates on the preparation will

be posted in the World-Wide Web page at http://www.astro.uiuc.edu/mcnews/MC IAUsymp.html
and at http://www.astro.uni-wuerzburg.de/~grebel/MC IAUsymp.html.

Please send comments and suggestions to mcnews@astro.uiuc.edu. Thank you.

The Magellanic Clouds Newsletter on the World-Wide Web

The Magellanic Clouds Newsletter has been available on the world-wide web for some time now (URLs see above). The present and past issues of the Newsletter can be downloaded either as L^{A}TEX files or as postscript files. The L^{A}TEX template is available as well. We update the compilation of meetings of potential interest for Magellanic Clouds researchers twice a month. A special web page is dedicated to a proposed IAU Symposium on the Magellanic Clouds (see above). Job advertisements of positions related to the Magellanic Clouds are posted as soon as we receive them. We provide links to other web sites related to Magellanic Clouds research and additional links of interest such as other newsletters, ground-based southern observatories, and space observatories. If you have suggestions for improvements, please let us know.

Abstracts of Refereed Papers

The Evolution of Ultraviolet Emission Lines From Circumstellar Material Surrounding SN 1987A

George Sonneborn^{1,2}, Claes Fransson^{3}, Peter Lundqvist^{3}, Angelo Cassatella^{4}, Roberto Gilmozzi^{5}, Robert P. Kirshner^{6}, Nino Panagia^{7,8}, and Willem Wamsteker^{9}

1 Laboratory for Astronomy and Solar Physics, Code 681, NASA/Goddard Space Flight Center, Greenbelt, MD 20771 2 Institut d’Astrophysique de Paris, 98 bis boulevard Arago, 75014 Paris, France 3 Stockholm Observatory, S-133 36 Saltsjöbaden, Sweden 4 Istituto di Astrofisica Spaziale, CNR, CP 67, I-00044, Frascati, Italy 5 European Southern Observatory, Karl-Schwarzschild Str. 2, 85748 Garching bei München, Germany 6 Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138 7 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218 8 Affiliated with the Astrophysics Division, Space Sciences Department of ESA 9 IUE Observatory, ESA-VILSPA, Casilla 50727, E-28080 Madrid, Spain

The presence of narrow high-temperature emission lines from nitrogen-rich gas close to SN 1987A has been a principal observational constraint on the evolutionary status of the supernova’s progenitor. A new analysis of the complete five-year set of low and high resolution IUE ultraviolet spectra of SN 1987A (1987.2–1992.3) provide fluxes for the N v 1240, N iv] 1486, He ii 1640, O iii] 1665, N iii] 1751, and C iii] 1908 lines with significantly reduced random and systematic errors and reveals significant short-term fluctuations in the light curves. The N v, N iv], and N iii] lines turn on sequentially over 15 to 20 days and show a progression from high to low ionization potential, implying an ionization gradient in the emitting region. The line emission turns on suddenly at 83 ± 4 days after the explosion, as defined by N iv]. The N iii] line reaches peak luminosity at 399 ± 15 days. A ring radius of (6.24 ± 0.20) × 10^{17} cm and inclination of 41.0 ± 3.9 degrees is derived from these times, assuming a circular ring. The probable role of resonant scattering in the N v light curve introduces systematic errors that leads us to exclude this line from the timing analysis. A new nebular analysis yields improved CNO abundance ratios of N/C= 6.1±1.1 and N/O= 1.7±0.5, confirming the nitrogen enrichment found in our previous paper. From the late-time behavior of the light curves we find that the emission originates from progressively lower density gas and that the emitting region has a multicomponent density structure. We estimate the emitting mass near maximum (∼ 400 days) to be ∼ 0.047 solar masses, assuming a filling factor of unity and an electron density of 2.6 × 10^{4}cm^{−3}. These results are discussed in the context of current models for the emission and hydrodynamics of the ring.

Accepted by: The Astrophysical Journal

For preprints, contact sonneborn@fornax.gsfc.nasa.gov

High velocity motions inside the Hii region N103 of the Large Magellanic Cloud

P. Ambrocio-Cruz^{1,2}, A. Laval^{1}, M. Marcelin^{1}, P. Amram^{1}

1 Observatoire de Marseille, 2, Place Le Verrier, 13248 Marseille CEDEX 4, France. 2 Instituto de Astronomía, UNAM, Apdo. 70-264, México, D.F., 04510, México.

The gas of the Hii region N103 of the Large Magellanic Clouds surrounds the double cluster NGC 1850.

We have observed the Hii region with a scanning Fabry-Perot interferometer at Hα and [Oiii]5007 wavelengths. The kinematics of this field show high velocity motions. We discuss their origin: Supernova explosion or particularly strong stellar winds.

By calculating the energy input inside the gas, we show that it is unlikely that the high velocity motions are due to the stellar winds of the embedded stars. The required number of O stars is too high for the known stellar content, and the extremely faint [Oiii]5007 emission gives a strong limit to the number of stars hotter than 37000 ^{o}K. The model of an old Supernova explosion agrees better to the energetic balance, if it occurred inside a cavity already driven by the stellar winds. Since the extrapolation of the IMF of the cluster leads to think that at least three massive stars have already exploded, such an assumption is the most plausible.

Then the nebula N103 is linked to two supernova remnants of different ages. The oldest one can be represented by a bubble, 152 pc wide, seen projected against the Hii region, and probably lying at the edge of the Hii region. The exciting stars of the nebula are actually members of the LMC cluster NGC 1850B; they provide a photon flux large enough to ionize the quiet part of the gas.

Accepted by: Astronomy and Astrophysics

For preprints, contact ambrocio@obmara.cnrs-mrs.fr

Ultraviolet Imaging Telescope Observations of OB Stars in the N 11 Region of the LMC

Joel Wm. Parker^{1,2}, Jesse K. Hill^{1}, Ralph C. Bohlin^{3}, Robert W. O’Connell^{4}, Susan G. Neff^{5}, Morton S. Roberts^{6}, Andrew M. Smith^{5}, and Theodore P. Stecher^{5}

1 Hughes STX Corporation, Code 681, GSFC/NASA, Greenbelt, MD 20771 2 current address: SwRI, Suite 426, 1050 Walnut Street, Boulder, CO 80302 3 Space Telescope Science Institute, Baltimore, MD 21218 4 Astronomy Department, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903 5 Laboratory for Astronomy and Solar Physics, Code 681,GSFC/NASA, Greenbelt, MD 20771 6 National Radio Astronomy Observatory, Edgemont Road, Charlottesville, VA 22903

We present an analysis of far-ultraviolet (FUV: 1300-1800 Å) and optical (U, B, and V) data of the stellar and nebular content of the OB associations LH 9, 10, and 13 in the Large Magellanic Cloud region N 11. The FUV images from The Ultraviolet Imaging Telescope strongly select the hot O and B stars; over 1900 stars were detected in the FUV to a limiting magnitude of m_{152} = 17 mag. The resulting FUV photometry combined with optical ground-based data indicate there are approximately 88 confirmed or candidate O stars in the LH 9, 10, and 13 fields alone (in an area of ∼ 41 arcmin^{2}), and possibly as many as 170 to 240 O-type stars within the entire 40 arcmin-diameter field of view.

Accepted by: The Astrophysical Journal

For preprints, contact joel@shasta.space.swri.edu
Also available by anonymous ftp at k2.space.swri.edu/pub/uit n11.ps
or from the URL http://k2.space.swri.edu/papers/papers.html

An Hi Aperture Synthesis Mosaic of the Small Magellanic Cloud

L. Staveley-Smith,^{1} R.J. Sault,^{1} D. Hatzidimitriou,^{2} M.J. Kesteven^{1} and D. McConnell^{1}

1Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 2121, Australia 2Physics Department, University of Crete, 714.09 Heraklion, Crete, Greece

We present the results of a survey of neutral hydrogen emission in the Small Magellanic Cloud (SMC) with the Australia Telescope Compact Array (ATCA). The survey consists of a mosaic of 320 separate pointings of the 375-m array, resulting in a resolution of 1.6 arcmin (28 pc, for a distance of 60 kpc) over a field of 20 deg^{2}. The rms brightness temperature sensitivity is 1.4 K, corresponding to an Hi column-density sensitivity of 4 × 10^{18} cm^{−2} for each velocity channel of width 1.6 kms^{−1}. The Hi distribution is complex and, on scales < 1 kpc, appears to be dominated by the effects of expanding Hi shells, which are probably driven by the combined effects of supernovae and stellar winds from massive stars. The picture of the SMC that arises from the current data seems to challenge the earlier belief that the SMC consists of two or more spatially separate structures with different systemic velocities. We find that the observed multiple components are, in many cases, caused by the combined effects of the numerous shells and supershells. Altogether, we identify six supershells (defined here as those with radii greater than 300 pc) and 495 giant shells. For each of these, we measure positions, radii, velocities and expansion rates, and derive ages and kinetic energy requirements. The apparent age distribution of shells is remarkably narrow, with a mean age of 5.4 Myr and an intrinsic dispersion of 2 Myr. Southern shells appear to be older, on average, by 2.5 Myr. The kinetic energy of the shells is a large fraction of the gravitational binding energy of the SMC, implying that further disintegration of the SMC will occur with time, and especially at the next close passage with the LMC or the Galaxy, unless the SMC possesses a massive halo. Because of their interferometric nature, the images presented here are insensitive to structures of size > 0.6^{∘}, and should not be used for deriving total Hi column densities. Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact lstavele@atnf.csiro.au
Also available by anonymous ftp at ftp://ftp.atnf.csiro.au/pub/data/smc/
or from the URL http://www.atnf.csiro.au/~lstavele

Ring nebula and bipolar outflows associated with the B1.5 supergiant Sher #25 in NGC 3603

Wolfgang Brandner^{1}, Eva K. Grebel^{1},^{2}, You-Hua Chu^{2}, Kerstin Weis^{2},^{3}

1 Astronomisches Institut der Universität Würzburg, Am Hubland, D-97074 Würzburg, Germany 2 Astronomy Department, University of Illinois at Urbana-Champaign, 1002 West Green St., Urbana, IL 61801, USA 3 Institut für Theoretische Astrophysik der Universität Heidelberg, D-69121 Heidelberg, Germany

We have identified a ring-shaped emission-line nebula and a possible bipolar outflow centered on the B1.5 supergiant Sher #25 in the Galactic giant Hii region NGC 3603 (distance 6 kpc). The clumpy ring around Sher #25 appears to be tilted by 64^{∘} against the plane of the sky. Its semi-major axis (position angle ≈ 165^{∘}) is 6.9′′ long, which corresponds to a ring diameter of 0.4 pc. The bipolar outflow filaments, presumably located above and below the ring plane on either side of Sher #25, show a separation of ≈ 0.5 pc from the central star.

High-resolution spectra show that the ring has a systemic velocity of V_{LSR} = +19 km s^{−1} and a de-projected expansion velocity of 20 km s^{−1}, and that one of the bipolar filaments has an outflow speed of ∼83 km s^{−1}. The spectra also show high [Nii]/Hα ratio, suggestive of strong N enrichment. Sher #25 must be an evolved blue supergiant (BSG) past the red supergiant (RSG) stage. We find that the ratio of equatorial to polar mass-loss rate during the red supergiant phase was ≈ 16. We discuss the results in the framework of RSG–BSG wind evolutionary models.

We compare Sher #25 to the progenitor of SN 1987 A, which it resembles in many aspects.

Submitted to: Astrophysical Journal Letters

For preprints, contact brandner@astro.uni-wuerzburg.de
Also available from the URL http://www.astro.uni-wuerzburg.de/supplements.html

Abstracts of Non-Refereed Papers

A photometric study of NGC 458

V. Ripepi^{1} and E. Brocato^{2}

1 Dipartimento di Fisica, UNICAL, Arcavacata di Rende, 87036 Rende (CS) Italy 2 Osservatorio Astronomico di Collurania, Via Mentore Maggini, 64100 Teramo, Italy

We present a CCD investigation of the poorly known young SMC globular cluster NGC 458. The NTT data presented here allowed us to study in detail the more internal regions of the cluster which are less contaminated by the field. On the basis of theoretical isochrones, a preliminary evaluation of the age is also given.

To appear in: Proceedings of the Elba Meeting “Stellar Ecology”, ed. R.T. Rood

For preprints, contact vincenzo@astr1pi.difi.unipi.it

Master Thesis Abstract

Age Determination For Two Globular Clusters in the Small Magellanic Cloud

Andrea Dieball^{1}

1 Sternwarte der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany

I analyzed B, V CCD photometry of the two globular clusters NGC 419 and NGC 416 in the eastern wing of the Small Magellanic Cloud (SMC). CCD photometry for both clusters was obtained with the 1.52 Danish telescope at ESO, La Silla. For NGC 416 I also used archival WFPC2 HST data. The ages of the clusters, of the surrounding field, and of an additional field in the wing of the SMC were determined by fitting isochrones from the Padua group (e.g., Bertelli et al. 1994, A&AS, 106, 275) to the colour-magnitude diagrams (CMDs). The CMDs show strong field star contamination. I therefore developed a program for statistical field star subtraction.

• For the cluster NGC 419 I derived an age of 1.3 ± 0.5 Gyr and a reddening of E_{B−V} = 0.03 mag with (m − M)_{0} = 18.9 mag.

• At least two pronounced young field populations could be discerned around NGC 419. The youngest field star population has an age of 12 to 50 Myr and a reddening of E_{B−V} = 0.01 mag. The second field population is somewhat older (130±30 Myr) and more strongly reddened (E_{B−V} = 0.06 mag), suggesting that it might be located behind the younger one. Isochrones with a metallicity of Z = 0.008 fitted both field star populations best.

• For NGC 416 best fits were obtained for an age of 2.6 ± 0.6 Gyr and E_{B−V} = 0.1 to 0.15 mag with (m − M)_{0} = 18.6 mag.

• Three young populations could be distinguished in the field in which NGC 416 is embedded: 8 ± 2 Myr, 50 ± 13 Myr, and 160 ± 30 Myr, all with E_{B−V} = 0.1 mag.

• For the field in the wing I detected three different populations, two of which have ages similar to the ones derived from the field surrounding the clusters: 25 ± 6 Myr, E_{B−V} = 0.13 mag 130 ± 30 Myr, E_{B−V} = 0.15 mag 2 ± 0.5 Gyr, E_{B−V} = 0.1 mag.

The results based on the ESO data are in good agreement with the results of Durand & Hardy (1984, ApJ, 283, 552) and of Gardiner & Hatzidimitriou (1992, MNRAS, 257, 195).

Since the HST data go down to fainter magnitudes than the ground-based data, and since they show a well-defined main-sequence turn-off for NGC 416, they allowed me to carry out a more accurate age determination. Based on the HST data I find an age of 6.3 ± 1.6 Gyr for a best-fitting isochrone with a reddening of E_{B−V} = 0.08 ± 0.01.

Diploma Thesis work carried out at the University of Bonn under direction of Prof. Klaas S. de Boer and Prof. Ulrich Mebold. Completed 1996 May. For preprints of forthcoming papers, contact adieball@astro.uni-bonn.de.

Job Advertisement

Postdoctoral Fellowship of the German Research Foundation at Bochum University

The Astronomical Institutes of the Universities of Bonn and Bochum are carrying out a joint research program on “The Magellanic Clouds and Other Dwarf Galaxies”. This project is financed by the German Research Foundation and comprises fellowships for PhD students and postdoctoral fellowships. Faculty members at the two universities involved in this project are Prof. K.S. de Boer, Prof. R.-J. Dettmar, Prof. J.V. Feitzinger, Dr. R. Hanuschik, Prof. U. Klein, Prof. U. Mebold, Dr. T. Richtler, Prof. W. Seggewiss, and Prof. Th. Schmidt-Kaler. For more information on the research

project, see http://aibn55.astro.uni-bonn.de:8000/~software/Dwarfs group.html and
http://www.astro.ruhr-uni-bochum.de/kam/grad.html.

A new postdoctoral fellowship within the framework of this research project at the Astronomical Institute of Bochum University will be available starting January 1, 1997. In accordance with remunerations of fellowships of the German Research Foundation, the tax-free salary for this position is DM 2400 (plus DM 300 for married postdocs).

Applications are invited from candidates with experience in structure and evolution of galaxies or in the physics of the interstellar medium. Please send your application documents and any inquiries to the speaker of the research group, Prof. U. Klein (e-mail: uklein@astro.uni-bonn.de; postal address: Radioastronomisches Institut der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany). The deadline for applications has been extended to October 31, 1996.

——–

We post job advertisements at the URLs

                  http://www.astro.uiuc.edu/mcnews/MCjobs.html and
                http://www.astro.uni-wuerzburg.de/~grebel/MCjobs.html

free of charge as soon as they become available.